OpenStudy (anonymous):

Differentiate function r(t)= 5t^(3/5)

OpenStudy (anonymous):

use formula f(x)= x^n therefore f'(x)=nx^(n-1)

OpenStudy (anonymous):

hint n=3/5

OpenStudy (anonymous):

cool I was a little confused there its 3t^(-2/5)

OpenStudy (anonymous):

good job

OpenStudy (anonymous):

\[5(3/5) t ^{3/5 - 1}\]